# TDAvec: Computing Vector Summaries of Persistence Diagrams for Topological Data Analysis in R and Python

### v1

@@ III-A Statement of need and Short Package Description
1. It expands the list of implemented vector summaries of PDs by providing vectorization of eight functional summaries found in the TDA literature: Betti function, persistence landscape function, persistence silhouette function, persistent entropy summary function [AGDST20], Euler characteristic curve [RW14], normalized life curve [CL22], persistence surface [AEK+17] and persistence block [CILS22] (see the Appendix for their definitions).
@@ III-B Usage Examples
    > devtools::install_github("ALuchinsky/tdavec")
After downloading the library, you can use functions such as computePL, computePS, etc., to calculate the corresponding vectorization summaries of persistence diagrams.
@@
Created point clouds can be converted into persistence diagrams using such functions as ‘ripsDiag‘ from ‘TDA‘ package. Each of the for each of the diagrams we can calculate Persistence Landscape summary with the help of ‘computePL(diagram, homDim, x)‘ function. In figure 2 you can see the result.
@@
The python interface of ‘TDAvec‘ library can be installed from the GitHub repository
@@
You can directly access all available in R interface unctions of the package, like ‘computePL‘, etc. In addition, python implementation follows the scikit-learn interface, so it uses a class-oriented approach. After installing the library, you can load all the objects with:
@@ Appendix A Definitions of the summary functions in TDAvec
3) The persistent entropy summary function:
@@
where $\beta_{0},\beta_{1},\ldots,\beta_{d}$ are the Betti curves corresponding to persistence diagrams $D_{0},D_{1},\ldots,D_{d}$ of dimensions $0,1,\ldots,d$ respectively, all computed from the same filtration.

### v2

@@ III-A Statement of need and Short Package Description
1. It expands the list of implemented vector summaries of PDs by providing vectorization of eight functional summaries found in the TDA literature: Betti function, persistence landscape function, persistence silhouette function, persistent entropy summary function [AGDST20], Euler characteristic curve [RW14], normalized life curve [CL22], persistence surface [AEK+17] and persistence block [CILS22]. In addition, in the updated version o the article such vectorizations as simple simple statistics, Tropical Coordinates, and Template Functions [PMK23] were added. In the Appendeix you an see definitions of theese vetorizations.
@@ III-B Usage Examples
    > devtools::install_github("uislambekov/TDAvec")
After downloading the library, you can use functions such as computePersistenceLandscape, computePersistenceSilhouette, etc., to calculate the corresponding vectorization summaries of persistence diagrams.
@@
Created point clouds can be converted into persistence diagrams using such functions as ripsDiag from TDA package. Each of the for each of the diagrams we can calculate Persistence Landscape summary with the help of computePerststenceLandscape(diagram, homDim, x) function [^3] Note that the names of the vectorization functions have changed during the last update. In figure 2 you can see the result.
@@
The python interface of TDAvec library can be installed from the GitHub repository
@@
You can directly access all available in R interface unctions of the package, like computePerststenceLandscape, etc. In addition, python implementation follows the scikit-learn interface, so it uses a class-oriented approach. After installing the library, you can load all the objects with:
@@ Appendix A Definitions of the summary functions in TDAvec
3) The persistent entropy summary function is defined in the following way:
@@
where $\beta_{0},\beta_{1},\ldots,\beta_{d}$ are the Betti curves corresponding to persistence diagrams $D_{0},D_{1},\ldots,D_{d}$ of dimensions $0,1,\ldots,d$ respectively, all computed from the same filtration. Note that unlike all other vectorizations the Euler characteristics curve is calculated using all available dimenstions.
@@
10) Simple statistics:
For a given persistence diagram $D=\{(b_{i},d_{i})\}_{i=1}^{N}$ (corresponding to a specified homological dimension), computeStats() calculates descriptive statistics of the birth, death, midpoint (the average of birth and death), and lifespan (death minus birth) values. Additionally, it computes the total number of points and entropy of the lifespan values. Points in $D$ with infinite death values are ignored.
11) Tropical Coordinates:
This vectorization for each homologial dimension computes the following seven tropical coordinates based on the lifespans (or persistence) $\lambda_{i}=d_{i}-b_{i}$:
  • $F_{1}=\max_{i}\lambda_{i}$.
  • $F_{2}=\max_{i<j}(\lambda_{i}+\lambda_{j})$.
  • $F_{3}=\max_{i<j<k}(\lambda_{i}+\lambda_{j}+\lambda_{k})$.
  • $F_{4}=\max_{i<j<k<l}(\lambda_{i}+\lambda_{j}+\lambda_{k}+\lambda_{l})$.
  • $F_{5}=\sum_{i}\lambda_{i}$.
  • $F_{6}=\sum_{i}\min(r\lambda_{i},b_{i})$, where $r$ is a positive integer.
  • $F_{7}=\sum_{j}\big{(}\max_{i}(\min(r\lambda_{i},b_{i})+\lambda_{i})-(\min(r\lambda_{j},b_{j})+\lambda_{j})\big{)}$.
    12) Template Function Vectorization
    For a given persistence diagram $D=\{(b_{i},d_{i})\}_{i=1}^{N}$ (corresponding to a specified homological dimension), this vectorization uses a collection of tent template functions defined by
    $G_{(b,p),\delta}(D)$ $=\sum_{i=1}^{N}\max\left\{0,1-\frac{1}{\delta}\max\left(|b_{i}-b|,|p_{i}-p|\right)\right\},$
    where $p_{i}=d_{i}-b_{i}$ (persistence), $b\geq 0$, $p>0$ and $0<\delta<p$. The point $(b,p)$ is referred to as the center. Points in $D$ with infinite death values are ignored.